OpenStudy (anonymous):

Solve this problem in your notebook using all four steps. Harvey is 3 times as old as Jane. The sum of their ages is 48 years. Find the age of each. Jane is________________a0 years old. Harvey is________________a1 years old.

OpenStudy (anonymous):

h + j = 48 h = 3j now sub 3j in for h 3j + j = 48 4j = 48 j = 12 h = 3j h = 3(12) h = 36 jane is 12 harvey is 36